A needle with an internal diameter of 0.38 mm and length 4.1 cm is used to deliver an intravenous fluid of density 1050 kg m−3 and viscosity 12×10−4 Pa s into a vein.

If the pressure difference across the needle is 8.0 kPa then how many hours will it take to empty 1 litre of IV fluid through the needle? (to 2 s.f)
1
Expert's answer
2020-04-14T09:39:38-0400

Assuming that the needle is of form of cylindrical pipe, according to Poiseuille law, the laminar flow through the needle is "Q = \frac{\pi d^4 \Delta p }{128 \eta l}", where:

"d" is the diameter of the needle, "\Delta p" is the pressure difference across the needle,

"\eta" is viscosity of the fluid, "l" is the length of the needle.


Substituting the values into Poiseuille formula, obtain "Q = 8.32 \cdot 10^{-8} \frac{m^3}{s}".

One litre is equal to "10^{-3} m^3", hence the time to empty this volume through the needle is "t = \frac{V}{Q} = \frac{10^{-3}m^3}{8.32 \cdot 10^{-8} \frac{m^3}{s}} = 12023.23 s \approx 3.34 h".